OpenStudy (anonymous):

There are 8 people at a picnic and their ages are as follows 15, 5, 2, 20, 7, 30, 40, 23. A group of 3 people is picked at random. a) How many groups are possible? b) How many groups are there with a combined age greater than 27?

OpenStudy (anonymous):

The order with which the people are picked doesn't seem to matter, so you use the combination formula: \[\large_{8}C_{3}=\frac{8!}{3!(8-3)!}=56\]

OpenStudy (anonymous):

I'm sure there's a quick way to figure out part (b), but so far I can think of it. Count up all the combinations of three of the numbers that sum to something greater than 27. \[2,~5,~7,~15,~20,~23,~30,~40\] 2, 5, 23 2, 5, 30 2, 5, 40 2, 7, 20 2, 7, 23 2, 7, 30 2, 7, 40 and so on.

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

I suppose a quick way is to count all the groups of people that already contain someone older than 27. We'd then have \[2\cdot\large_{7}C_{2}=2\cdot21=42\] because any group containing the 30 or 40 year old will automatically qualify as "greater than 27." Add up the remaining combinations that qualify.